### 23 Jan 2024 v1

# DALex: Lexicase-like Selection via Diverse Aggregation

Abstract
Lexicase selection has been shown to provide advantages over other selection algorithms in several areas of evolutionary computation and machine learning. In its standard form, lexicase selection filters a population or other collection on the basis of randomly ordered training cases that are considered one at a time. This iterated filtering process can be time-consuming, particularly in settings with large numbers of training cases. Such settings include many symbolic regression and deep learning applications. In this paper, we propose a new method that is nearly equivalent to lexicase selection in terms of the individuals that it selects, but which does so significantly more quickly. The new method, called DALex (for Diversely Aggregated Lexicase), selects the best individual with respect to a weighted sum of training case errors, where the weights are randomly sampled. This allows us to formulate the core computation required for selection as matrix multiplication instead of recursive loops of comparisons, which in turn allows us to take advantage of optimized and parallel algorithms designed for matrix multiplication for speedup. Furthermore, we show that we can interpolate between lexicase-like behavior and the behavior of “relaxed” variants of lexicase selection, such as epsilon lexicase selection and batch lexicase selection, by adjusting a single hyperparameter, named “particularity pressure,” which represents the importance granted to each individual training case. Results on program synthesis, deep learning, symbolic regression, and learning classifier systems demonstrate that DALex achieves significant speedups over lexicase selection and its relaxed variants while maintaining almost identical problem-solving performance. Under a fixed computational budget, these savings free up resources that can be directed towards increasing population size or the number of generations, enabling the potential for solving more difficult problems.
Keywords:
Lexicase Selection Learning Classifier Systems Genetic Programming Symbolic Regression Deep Learning
1 Introduction
Genetic algorithms [27] is a subfield of Evolutionary Computation that uses concepts from biological evolution—random variation and survival of the fittest—to solve a wide array of difficult problems. Genetic programming (GP) is a subfield of genetic algorithms that evolves programs or functions that take some inputs and produce some outputs. Common subfields of GP include software synthesis, in which a population of programs evolves to satisfy user-defined training cases, and symbolic regression (SR), in which a population of mathematical expressions evolves to fit a regression dataset. Central to these genetic algorithms is the concept of parent selection, in which members of the current population with desirable characteristics are chosen as the starting point from which to create the next generation of individuals. Lexicase selection [25] and epsilon lexicase selection [33] are state-of-the-art selection algorithms developed for the discrete-error domain and the continuous-error domain, respectively. The key idea behind lexicase selection is that it can be helpful to disaggregate the fitness function, selecting parents by considering training cases one at a time in a random order, all the while filtering out individuals that are not elite relative to the other remaining individuals on the current training case. This selection method has been shown to maintain beneficial diversity in evolved populations [1, 19], especially in terms of “specialists”: individuals who may have a high total error but have a very low error on a subset of the training cases [21, 22]. However, due to its iterative nature, lexicase selection can often take a long time and there is not an obvious way to take advantage of parallelism or single-instruction-multiple-data (SIMD) architectures to speed up its runtime.
In this work, we reexamine the lexicase intuition behind deaggregating the fitness function and we develop an efficient selection method based on randomly aggregating the fitness function at each unbatched selection event. Specifically, we quantify the idea that training cases that occur earlier in a lexicase ordering exert greater selection pressure by taking a randomly weighted average of each individual’s error vector. We formulate this selection mechanism in terms of matrix multiplications, which allows us to take advantage of modern advancements in matrix multiplication algorithms and SIMD architectures to achieve significantly faster runtime compared to lexicase selection. In this work, we distinguish between an unbatched selection event, which takes in the error vectors of the population and return the index of the selected parent, and a batched selection event, which takes in the error vectors of the population and returns an equal number of indices representing the selected parents. For a population of size $n$, creating the next generation would take $n$ unbatched selection events or 1 batched selection event. While we present the theory in terms of an unbatched selection event to simplify the discussion, we implement our selection method and present results using a batched selection event.
This paper is organized as follows: In Section 2, we give a brief overview of lexicase selection and its variants. In Section 3, we describe the Diversely Aggregated Lexicase (DALex) selection method and give a brief theoretical treatment of its properties. In Section 4, we describe our experiments and results in three popular domains in which lexicase selection has had demonstrated success. Empirical results show that DALex replicates the problem-solving success of lexicase selection and its most successful variants while offering significantly reduced runtime.
2 Background and Related Work
Lexicase selection is a parent selection method that assesses individuals based on each training case in turn, instead of constructing a single scalar fitness value for each individual [25]. During each selection event, the training cases are randomly shuffled. For each training case in the order determined by the random shuffle, candidate individuals in the population that are not “elite” with respect to that training case, i.e., have an error greater than the minimum error on that training case among the remaining candidates, are filtered out. If at any point only a single individual remains, that individual is selected. If all of the training cases are exhausted, then a random individual is selected from those remaining. In other words, lexicase selection chooses an individual based on the first training case, using the rest of the cases in order to break ties. This mechanism has been shown to better maintain population diversity than methods based on aggregated fitness measures [19], which has been shown to improve problem-solving performance [17]. Lexicase selection and its variants have been successfully applied in many diverse problem domains beyond software synthesis, such as learning classifier systems [3], symbolic regression [33], SAT solvers [36], deep learning [11], and evolutionary robotics [44].
In continuous-valued domains like SR, it is often unlikely for more than one individual to share the minimum error on a training case. Therefore, the selection pressure exerted by each training case will be too large, and lexicase selection will not proceed beyond one training case. To combat this, the epsilon-lexicase selection method was developed as a relaxation of lexicase selection [33]. Instead of requiring individuals to have the minimum error out of the current candidates on a training case, epsilon-lexicase filters out individuals with errors exceeding an epsilon threshold above the minimum error. This epsilon threshold is adaptively determined by the population dynamics, specifically the median of the absolute deviations from the median error on each training case.
Batch-lexicase [3, 35, 40] is another lexicase relaxation with improved performance on noisy datasets. Instead of considering each training case in turn, batch-lexicase considers groups of training cases, filtering out individuals whose average accuracy on that group is lower than some threshold. Batch-lexicase has been successfully applied to the field of Learning Classifier Systems (LCS) in Aenugu et al. [3], showing improved performance on noisy datasets and improved generalization with limited training data compared to lexicase selection.
There have been many attempts to speed up lexicase selection. Dolson et al. [12] showed that the exact calculation of lexicase selection probabilities is NP-hard. However, Ding et al. [10] are able to calculate an approximate probability distribution from which to sample individuals. They show that their selection method achieves significantly faster runtime while achieving similar selection frequencies compared to lexicase selection. Ding et al. [9] also propose to use a weighted shuffle of training cases so that more difficult cases are considered first. They show that this technique can significantly reduce the number of training cases considered while suffering minimal degradation of problem-solving power. Batch tournament selection (BTS) [35] orders the training cases by difficulty, then groups them into batches. For each batch of cases, BTS then performs tournament selection once, using the average error on that batch as fitness. Compared to epsilon-lexicase selection, they show competitive problem-solving performance on SR datasets with significantly faster runtime.
Furthermore, many selection methods based on aggregated fitness measures have been proposed for multiobjective genetic algorithms. Fitness sharing is a selection method that attempts to maintain diversity by penalizing individuals for being too close to other individuals in the population, with respect to a user-defined distance metric [8]. On the other hand, implicit fitness sharing does not require a distance metric, and instead scales the reward for each training case by the rest of the population’s performance on that training case before computing the aggregated fitness function [34]. Historically assessed hardness (HAH) is another form of fitness sharing in which errors on each training case are scaled by the population’s historical success rate on that problem [29]. In addition to using a distance metric, NSGA-II sorts individuals into pareto-fronts and conducts tournament selection using the nondomination rank and local crowding factor as fitness values [7].
This work lies at the intersection of the lexicase variants mentioned above but takes a different direction compared to the speedup methods developed so far. For each unbatched selection event, we sample a random weighting of training cases. Then, for each individual, we compute a weighted average of the training case errors and select the individual with the lowest error with respect to the weighted average. In contrast to implicit fitness sharing (IFS) or HAH, we use weights that are randomly chosen, span many more orders of magnitude, and are resampled at each selection event instead of at each generation. In contrast to NSGA-II, this method only selects non-dominated individuals because all of the training case weights are positive after the softmax operation. However, like epsilon and batch lexicase, and in contrast to lexicase selection, this method may select individuals outside of the “Pareto boundary” as defined in La Cava et al. [31]. We formulate this selection method as a vectorized matrix multiplication, selecting all individuals in one go. This allows us to take advantage of modern SIMD architecture and algorithmic advances in matrix multiplication to achieve significantly faster runtime compared to lexicase selection.
3 Diversely Aggregated Lexicase
3.1 Description
Diversely Aggregated Lexicase selection (DALex) operates by selecting the individual with the lowest average error with respect to a randomly sampled set of weights. It is parameterized by the shape and scale of the distribution from which these weights are sampled. In most of our experiments, we fix the shape to be normal and vary the scale by changing the standard deviation of the distribution. We start with a population of $n$ individuals, where each individual is evaluated on $m$ training cases. We sample an importance score for each training case from the distribution $\mathcal{N}(0,{\tt std})$ where the standard deviation std is a tunable hyperparameter which we call the particularity pressure. Generally speaking, a training case’s importance score quantifies how many times more important it is compared to the rest of the training cases. To obtain the training case weights, we apply the softmax function to the importance scores. This allows us to turn differences between the values of importance scores into differences between the magnitudes of the training case weights. For each individual, we compute its average error w.r.t. the weighting as the dot product of the weight vector and the individual’s error vector. Finally, we select the individual with the lowest average error.
In practice, we conduct all selection events at once, batching multiple weight vectors into a weight matrix and using matrix multiplication in place of the dot product. We also perform an initial “pre-selection” step that differs slightly from other lexicase implementations due to the batched nature of our method: we group individuals into equivalence classes based on their error vectors, select equivalence classes using DALex, and then choose a random individual from each selected class [18].
For a population of $n$ individuals evaluated on $m$ training cases, there will be $k\leq n$ distinct equivalence classes, so the algorithm multiplies a $k\times m$ error matrix with a transposed $n\times m$ weight matrix, giving a worst-case asymptotic runtime of $O(m^{2}n)$, which is the same as that of lexicase selection. However, due to advances in the theory of matrix multiplication, the runtime of DALex can be reduced to $O(n^{\omega(\log_{n}m)})$ using methods such as the Coppersmith-Winograd algorithm [6].
3.2 Motivation
For a better understanding of DALex, we provide two intuitions linking it to lexicase selection.
First, we argue that DALex prompts us to reconsider a central tenet of lexicase selection. The success of lexicase selection is commonly attributed to its deaggregation of the fitness function. Aggregated selection methods like tournament or fitness-proportional selection, the reasoning goes, are unable to maintain effective problem solving diversity the same way lexicase selection does. To rationalize our results on DALex, we instead propose the slightly different view that the discrepancy is due to the single unweighted aggregation event preceding these selection methods, which causes a loss of information. Because lexicase selection utilizes the entire high-dimensional error vector instead of simply its mean or sum, it is able to more accurately select promising individuals. Since there are many diverse aggregation events in DALex, it is able to fully utilize the information contained in the individuals’ error vectors, and is therefore able to replicate lexicase selection’s success. Even though each individual selection event operates on a single-dimensional total error, diversity arises from different random weights promoting different individuals in different selection events.
Second, we draw a correspondence between the random orderings in lexicase selection and the random weightings in DALex. Given a random ordering of training cases in lexicase selection, the training cases occurring earlier in the ordering are more relevant to what individual is chosen at that selection event, i.e. exert higher selection pressure. The first training case is paramount, and only if multiple individuals are equally good at the first case do we consider the second case, and so on. In fact, in GP problems, the last case in the ordering is often irrelevant because an individual is usually selected after considering a fraction of the total number of cases. In DALex, the importance scores or training case weights assign explicit values to this notion of selection pressure. Given training case weights of $[1.0,~{}0.01,~{}0.1]$ for example, the first case would be ten times as important as the third case, which in turn is ten times as important as the second case. In the limit of infinite particularity pressure, the differences in importance scores tend towards infinity, so the difference in magnitudes of training case weights tends to infinity, and we recover lexicase selection. As the number of training cases increases or the number of orders of magnitude spanned by the errors on each case increases, the particularity pressure needed to replicate lexicase selection increases. Empirically, we are able to replicate lexicase selection using modest particularity pressures such as 20.
One important result of this interpretation is that it explains why we call the standard deviation hyperparameter the particularity pressure. When importance weights are sampled with high standard deviation, the most important training case will be much more important than the second most, and so on, just like in lexicase selection. As the standard deviation decreases, the importance of a single training case decreases. In other words, high particularity pressure corresponds to increasingly lexicase-like selection [42]. We find that we can achieve similar successes to relaxed forms of lexicase selection such as batch-lexicase or epsilon-lexicase simply by choosing an appropriate particularity pressure. In this sense, DALex unifies the diverse selection methods lexicase, epsilon-lexicase, and batch-lexicase into a single selection method. It is important to note that DALex’s only theoretical guarantee is that of asymptotically lexicase-like selection with infinite particularity pressure. As decreasing the particularity pressure is a different way of relaxation compared to those proposed for epsilon-lexicase and batch-lexicase, we do not expect to be able to replicate the selection dynamics of these lexicase variants to an arbitrary degree of accuracy. However, empirical results show that varying the particularity pressure gives enough flexibility to replicate the successes of these relaxed lexicase variants on the problem domains for which they were designed.
3.3 Modifications
For domains suited for a relaxed lexicase variant such as epsilon-lexicase or batch-lexicase, we first standardize the population errors on each training case to have a mean of 0 and a variance of 1 before multiplying by the training case weights. While not necessary, we find that this normalization step helps DALex perform well across many problems with a single hyperparameter setting.
Of the problem domains we study, Learning Classifier Systems (LCS) [46] has a significantly different structure and therefore requires further modification. In LCS, where an individual represents a rule that matches a subset of the training cases, each individual will only have errors defined on a subset of the training cases. For these problems, we let the individual have error 0 on cases for which it is undefined. We also use the individual’s support vector, which has a 1 for each training case on which the individual is defined, and a 0 otherwise. To compute the individual’s average error, we take the dot product of the error vector with the weight vector and then normalize by dividing this value by the dot product of the support vector with the weight vector.
Algorithm 1: Diversely Aggregated Lexicase, batched selection
Data:
Output:
if Relaxed:
  $E$leftarrow $\frac{E-{\tt mean}(E,{\tt axis}=0)}{{\tt std}(E,{\tt axis}=0)}$ The standardized error matrix
end if
[…]
$W$leftarrow ${\tt softmax}(I,{\tt axis}=1)$ The training case weights
$F$leftarrow $\frac{EW^{T}}{SW^{T}}$ The normalized, weighted, average error for each individual
${\tt idx}$leftarrow ${\tt argmin}(F,{\tt axis=0})$ lowest error individuals w.r.t the weights $W$
return idx
In problems where individuals are defined on all training cases, the support is 1 everywhere, so the normalization constant $SW^{T}$ is simply the sum of the training case weights, which is 1 due to the softmax operation. Under this condition, our algorithm degenerates to the simpler case considered above. The complete algorithm pseudocode, with the augmentations described above to accommodate for LCS and SR problems, is described in Algorithm 1.
4 Experiments and Results
4.1 CBGP
We first compare DALex to lexicase selection on program synthesis problems, the class of problem for which lexicase selection was first developed. For these experiments, we use a particularity pressure of 20 As DALex can exactly simulate lexicase selection in the limit of infinite particularity pressure, we also assess the ability of DALex to replicate the lexicase selection probability distribution with a moderate particularity pressure. To situate our results with respect to other lexicase approximations we also compare to plexicase selection [^1] Fixed a bug in the downsampling implementation in the released version of [10][10]. For plexicase selection, we use the hyperparameter setting $\alpha=1$. We use the Code-Building Genetic Programming (CBGP) system developed by Helmuth et al. [37] and test on a suite of problems from the PSB1 Benchmark [23] on which lexicase selection has shown good performance. We follow the default settings in CBGP, evolving 1000 individuals for 300 generations and using UMAD [20] with a rate of 0.09 as the sole variation operator. We present results in both the full data and downsampled paradigms. In the full data paradigm, all individuals are evaluated on all training cases. In the downsampled paradigm, a subset of the training cases is sampled at each generation and used to evaluate individuals in that generation. Downsampling has been shown to benefit the performance of Genetic Programming as we can increase the number of generations and/or individuals in the population for a given total computational budget [24, 2, 26]. Additional downsampling methods utilizing information from the population to choose the downsampled cases have been proposed [4], but we only study the randomly downsampled paradigm. In our experiments, we use a downsampling rate of 25%. The specific problems we study are Compare String Lengths (CSL), Median, Number IO, Replace Space with Newline (RSWN), Smallest, Vector Average, and Negative to Zero (NTZ).
Following the recommendation of the PSB1 benchmark, we report problem-solving performance as number of successful runs out of 100. We define a successful run as one which produces an individual with zero error on both the training cases and the unseen testing cases.
Table 1: Success rates of GP runs on six benchmark problems in the full data paradigm. The success rates across the board are very similar, and no results were significantly different between the selection methods. We perform chi-squared tests and underline results that were significantly worse than lexicase with ($p<0.05$). No results were significantly better than lexicase.
Problem Success Rate Downsampled Success Rate
Lexicase DALex Plexicase ($\alpha=1$) Lexicase DALex Plexicase ($\alpha=1$)
CSL 0 0 0 0 0 0
Median 91 91 83 100 100 60
Number IO 99 99 100 100 100 100
RSWN 12 15 6 66 68 50
Smallest 100 100 100 100 100 100
Vector Average 100 100 99 100 100 100
NTZ 78 83 80 99 100 100
Table 1 compares the success rates of DALex, plexicase, and lexicase selection on six benchmark problems under the full and downsampled paradigms. No results were significantly different between the selection methods, and DALex achieves very similar success rates to lexicase selection. We find that DALex runs faster than lexicase selection but slower than plexicase selection. Detailed results can be found in the appendix.
To compare the selection dynamics of the three methods, we solve a GP problem using lexicase selection, and at each selection event we sample 50,000 individuals using each selection method. From this we build up the empirical probability distributions $\{p_{it}\}$ the probability of selecting individual $i$ at generation $t$ via lexicase selection, $\{p_{it}^{\prime}\}$ the probability of selecting individual $i$ at generation $t$ via DALex, and $\{p_{it}^{\prime\prime}\}$ the probability of selecting individual $i$ at generation $t$ via plexicase. To quantify the differences in probability distributions, we use the Jensen-Shannon divergence metric
$D_{t}=\frac{1}{2}\left[\sum_{i}p_{it}\log\left(\frac{2p_{it}}{p_{it}+q_{it}}\right)+\sum_{i}q_{it}\log\left(\frac{2q_{it}}{p_{it}+q_{it}}\right)\right]$
Where $q_{it}$ is $p^{\prime}_{it}$ or $p^{\prime\prime}_{it}$. Furthermore, for each run in which lexicase selection produces a generalizing individual, we track that individual’s lineage to find its ancestor $a_{t}$ in each generation $t$, of which there is only one per generation since we only use the single-parent variation operator UMAD. For each of these individuals we compute the probability ratio $r_{t}=\frac{q_{a_{t}t}}{p_{a_{t}t}}$ quantifying how much more likely DALex is to select the successful lineage than lexicase selection. We report the average of these two metrics over all generations and all runs.
(a) Probability Ratio
(b) JS-Divergence
Figure 1: Fidelity of the DALex approximation to lexicase selection on CBGP problems. Error bars show the bootstrapped 95% confidence intervals. The probability ratio of DALex selecting the successful lineage is very close to 1 in all problems except for Vector Average. The Jensen-Shannon divergence of the DALex selection distribution from the lexicase selection distribution is close to 0 in all problems except for Vector Average and Negative To Zero.
Figure 1 shows the results for these two metrics on the problems studied. In almost all problems, the probability of selecting the successful lineage is almost the same between DALex and lexicase. In contrast, the probability ratio of plexicase and lexicase deviates much more from 1 on many of the problems, indicating a less faithful approximation. Furthermore, the Jensen-Shannon divergences between the DALex and lexicase selection probability distributions are all very close to zero and much smaller than the divergences of the plexicase selection distribution from the lexicase selection distribution, indicating that DALex gives a close approximation of the lexicase selection dynamics. Finally, we expect increasing the value of the standard deviation to improve the fidelity of the DALex approximation.
As an ablation experiment, we examine the effect of the distribution from which importance scores are chosen on the problem-solving ability of DALex. We compare our default selection from a normal distribution to selection from a uniform distribution and shuffling a range of evenly spaced numbers. In each case the importance scores are scaled so that they have a standard deviation of 20 prior to softmaxing.
Table 2: Success rates of GP runs using different distributions for DALex. All distributions were scaled to a standard deviation of 20 unless otherwise noted. We perform chi-squared tests and underline results that were significantly worse than the normal distribution with ($p<0.05$). No results were significantly better than the normal distribution.
Problem Success Rate
 Normal  Uniform  Range
CSL 0 0 0
Median 91 88 87
Number IO 99 99 100
RSWN 15 15 18
Smallest 100 100 100
Vector Average 100 100 100
NTZ 83 65 66
NTZ (std=100) 79 85 72
The results are shown in Table 2. The different distribution choices perform almost identically on the problems considered except for the Negative To Zero problem. We hypothesize that this problem is more taxing for our lexicase approximations, and highlights the differences between the distributions used for DALex. Of the three distributions, the normal distribution has the largest kurtosis. Therefore, the largest importance scores sampled from the normal distribution will be farther apart than for the other two distributions. Since lexicase selection often selects an individual after considering only a fraction of the training cases [22], it is these largest importance scores which determine the success of DALex. The uniform distribution and random shuffle, on the other hand, lower kurtosis, and perform worse compared to the normal distribution. However, increasing the standard deviation recovers the success rate of the normal distribution.
4.2 Image Classification
To test the ability of DALex to mimic lexicase selection even with many training cases, we compare using DALex versus lexicase selection as the selection method in gradient lexicase. We use the popular VGG16 [39] and ResNet18 [16] neural networks on the CIFAR10 [30] dataset. This dataset is comprised of 32x32 bit RGB images distributed evenly across 10 classes. There are 50,000 training instances and 10,000 test instances. As shown in Ding et al. [9], lexicase selection often uses up to the entire training set to perform selection, especially towards the end of the genetic algorithm.
Gradient lexicase [11] is a hybrid of lexicase selection with deep learning, in which a population of deep neural networks (DNN) is evolved to fit an image classification dataset. In gradient lexicase, stochastic gradient descent on a subset of the training cases is used as the mutation operator. Lexicase selection proceeds by evaluating the population of DNNs on one training case at a time and keeping only the DNNs that correctly predict the training case’s label. In contrast, DALex evaluates each individual on the entire training set, assigns to each training case an error value of 0 if the case was correctly predicted or 1 otherwise, and computes a weighted sum of the error values. For this domain, we increase the particularity pressure tenfold to 200.
Since DALex is appropriate for both discrete-valued and continuous-valued errors, we also investigate the possibility of using cross-entropy loss instead of accuracy as the basis of selection.
We use a population size of $p=4$ and run evolution for a total of $200(p+1)$ epochs. For a more detailed description of other hyperparameters such as the SGD learning rate schedule, see Ding et al. [11].
[…]
Table 3: Performance of five selection methods on the CIFAR10 dataset using the VGG16 architecture. DALex acheives similar test accuracy and lower runtime compared to lexicase selection.
Selector VGG16 ResNet18
Test Accuracy Runtime (h) Test Accuracy Runtime (h)
Vanilla SGD $92.6\pm 0.1$ $1.0\pm 0.001$ $93.72\pm 0.03$ $1.0\pm 0.006$
Lexicase $93.7\pm 0.05$ $13.1\pm 0.5$ $95.0\pm 0.1$ $35\pm 1$
DALex accuracy (std=200) $93.8\pm 0.1$ $8.27\pm 0.05$ $95.2\pm 0.1$ $9.0\pm 0.1$
DALex losses (std=200) $93.9\pm 0.1$ $8.25\pm 0.01$ $95.2\pm 0.2$ $8.9\pm 0.1$
We find that DALex performs just as well as lexicase selection whether selecting on cross-entropy losses or accuracy, and additionally runs much faster. This is likely due to lexicase selection’s evaluation of individuals on each training case with a batch size of 1. The small batch size causes too much communication overhead between the cpu and gpu, resulting in a slower runtime. Furthermore, as training progresses and the neural nets approach 100% accuracy on the training set, lexicase selection begins to exhibit its worst-case runtime. As the ResNet18 model reaches higher training set accuracy earlier than the VGG16 model, lexicase selection takes more than twice as long to run. With DALex, however, both models have a consistent runtime that is much lower than the runtime of lexicase selection.
4.3 SRBench
[…]
The statistics we examine in this paper, following the SRBench defaults, are test $R^{2}$, training time, and model complexity. We report the median of these statistics over all 20 problems, using bootstrapping to obtain a 95% confidence interval.
[…]
(b) Runtime
(c) Model Size
[…]
4.4 Learning Classifier Systems
Finally, to show that DALex can reproduce the success of batch-lexicase, we use the learning classifier system (LCS) from Aenugu et al. [3] We choose the led24 dataset [5] for the stark contrast between the performances of lexicase selection and batch-lexicase selection. The led24 problem [5], taken from the UCI ML Repository, tests the ability of selection methods to cope with noise. This dataset, consisting of 3200 instances, contains 7 boolean attributes corresponding to whether each of 7 LED lights in a display is lit up. It has 10 categories, the digits 0 through 9, corresponding to which digit is shown on the LED display. In this dataset, each boolean attribute has a 0.1 probability having its value inverted. The evolved rule distribution for this problem is sparse, with most rules representing fewer than 5 data instances [9]. This property makes the led24 problem unsuitable for lexicase selection and motivates the development of batch-lexicase selection. Since batched lexicase selection is also a relaxed version of lexicase selection, just like epsilon lexicase, for this domain we also need to determine the particularity pressure via a hyperparameter search, and choose the value of 20.
As in Aenugu et al. [3], we conduct experiments in both the full data and partial data paradigms. In the full data paradigm, we use the entire led24 dataset, leaving out 30% for testing. In the partial data paradigm, we randomly downample the dataset, removing 40% of the instances. We repeat this downsample for each run, meaning the instances available for each run will be different. We then split the remaining 60% of the instances into a 40% training set and a 20% testing set.
(a) Accuracy (Full)
(b) Runtime (Full)
(c) Rulecount (Full)
(d) Accuracy (Partial)
(e) Runtime (Partial)
(f) Rulecount (Partial)
Figure 3: Performance of four selection methods on the led24 problem. Bootstrapped 95% confidence intervals are indicated by error bars in the bar plots and shading in the line plots. DALex equals the highest test accuracy while having very low runtime and generating the fewest rules out of the four methods.
Figure 3 displays the results of our experiments. In both domains, we find that DALex has both the highest test accuracy of the four selection methods and generates the least number of distinct rules, which suggests that it may have better generalization ability. Finally, DALex has a significantly lower runtime in both domains than lexicase and batch-lexicase selection.
5 Conclusion and Future Work
In this work, we presented the novel parent selection method DALex, which efficiently approximates the lexicase selection mechanism by using random aggregations of error vectors. The proposed method, parametrized by the particularity pressure, is flexible enough to replicate the successes of lexicase selection and its relaxed variants on multiple problem domains, and is robust to very large datasets. Furthermore, it has a consistent and low runtime, which frees up resources that can be used for the other aspects of evolution such as number of generations. We make the code used in these experiments available at https://github.com/andrewni420/DALex.
While we have used reasonable presets for the standard deviation hyperparameter in DALex without performing significant hyperparameter search, we do not anticipate that a single hyperparameter setting will work for all cases or even for all generations during the course of a GP run. Therefore, future work could adaptively determine the optimal standard deviation during the course of evolution using hyperparameter search or even a more radical paradigm such as autoconstruction [41, 43], Additionally, since DALex computes a randomly weighted fitness function, it can be combined with selection methods other than pure elitist selection, such as tournament or fitness-proportional selection [13] to increase diversity. Furthermore, the weighted aggregation formulation of DALex is not confined to parent selection algorithms. This randomly weighted averaging concept could potentially be used to transform the losses for each training case in deep learning. By introducing more stochasticity into training, this could enable models to be trained with large batch sizes without the performance degradation that comes from lack of stochasticity [28]. Alternatively, DALex could be used to randomly weight each node in the policy head of a reinforcement learner when sampling trajectories, improving the exploration of off-policy learning algorithms such as Double Q-learning [15] or SAC [14] and potentially even removing the need for entropy regularization in policy-gradient algorithms.
Appendix 0.A Program Synthesis Runtime Comparisons
During our program synthesis experiments, we record the time taken to select the parents of the next generation. We then take the average of this metric over all generations and all runs and report them in Figure 4. DALex takes up to 5x less time to run than lexicase selection. While having a similar or slower runtime compared to plexicase selection, DALex has shown better performance as presented in Table 1, where plexicase selection occasionally produces significantly worse results than lexicase selection.
Figure 4: Runtime comparison of three selecion methods on program synthesis problems. 95% confidence intervals obtained by bootstrapping.